THE B̃1/2 (^{2}Σ^{+}) − X̃^{2}Σ^{+} ELECTRONIC TRANSITIONS OF LaNH AND LaND

DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055; ANTHONY J. MERER, Department of Chemistry, University of British Columbia, Vancouver, BC, Canada V6T 1Z1; \underline{SCOTT J. RIXON}, Department of Physics and Astronomy, University of British Columbia, 6224 Agricultural Road, Vancouver, BC, Canada V6T 1Z1; THOMAS D. VAR-BERG, Department of Chemistry, Macalester College, St. Paul, MN 55105.

New high resolution spectra of the B̃1/2 (^{2}Σ^{+}) − X̃^{2}Σ^{+} transitions of LaNH and LaND, near 16000 cm^{−1}, have been obtained under jet-cooled conditions, following the reaction of laser-ablated lanthanum metal with NH_{3} and ND_{3}. Rotational analyses have given the ground state structure as r_{0}(La-N) = 1.932072(11) Å; r_{0}(N-H) = 1.01357(31) Å. Strong vibronic and spin-orbit coupling is found to occur between the B̃ state and the Ã^{2}Π state near 12600 cm^{−1}. This causes various interesting effects: the hyperfine line strengths in the B̃ − X̃ transition are unusual, showing that, though it is nominally ^{2}Σ − ^{2}Σ, both parallel and perpendicular transition moments are present; also the bending frequency of the B̃ state is raised by 150 cm^{−1} compared to the ground state, while its vibrational structure shows large l-dependent “spin-vibration” splittings which can be represented by A_{sv} l, where A_{sv} ≃ −14 cm^{−1}.